### v1

@@ 5.2 Pre-collapse structure
Figure 10: Properties of the 20 $M_{\odot}$ GENEC model before and then during a silicon/oxygen shell merger event occurring very shortly before collapse. The time gap between the two models shown is 15 s. On the left we show the convective Mach number, i.e., the ratio between the velocity predicted by mixing length theory over the sound speed and also the energy produced by nuclear burning. On the right is shown the density and entropy structure of the shell in question.
We know that shell mergers and enhanced mixing/burning prior to core collapse can have a strong impact on the stellar structure before collapse and thus impact the supernova explosion and evolution (Müller et al. 2016; Jones et al. 2017; Rizzuti et al. 2022). We can see this also in Fig. 10. In this figure, we display the convective Mach number, entropy and density profiles of the 20 $M_{\odot}$ GENEC model before and then during a silicon/oxygen shell merger event. These profiles are taken roughly 404040\,40s before collapse, and there is a time difference of 15 s between the two models shown. We see that during the shell merger there is a rapid amplification of the energy generation by nuclear burning, an increase by more than an order of magnitude in the entire shell. This is accompanied by a steep increase of the convective mixing as the convective Mach number increases by a factor 3 in the merged shell. Furthermore, this merger event shifts the shell interface. The shift is only by $0.01M_{\odot}$ however we note that the compactness, $\xi_{r}=\frac{M_{r}[M_{\odot}]}{(r[1000\rm km])}$, at the shell interface changes here by $8\%$ going from 0.427 to 0.464. These structural changes will impact the explosion dynamics. Therefore, knowing if the presence of these shell mergers before collapse are physical and are likely to happen is crucial. In this grid of models, shell mergers seem to be dependent on the stellar evolution code and the chosen prescriptions. These events are of course highly sensitive to differences in the network implementation and convective treatment, as such, exploration in further detail on the trigger of these shell mergers and their impact on the explosion is required.
Figure 11: Space-time diagrams, or Kippenhahn diagrams, of the $15M_{\odot}$ model for GENEC, left, and MESA, right. The grey zones show convective regions. The plot background shade displays the total energy input, the nuclear energy subtracted by the loss of thermal neutrinos in log scale. Blue represents negative, and red positive, energy contributions. On the right axis is shown the central abundance of a given selection of species to help identify the various burning phases. The black dashed line delimits the total mass of the model.
@@ 5.3 Explodability conditions
Figure 12: Evolution prior to collapse of different instances of the compactness parameter, $\xi_{Mr=2.5}$, $\xi_{M_{r}=M_{4}}$, see Eq. (23) for definition.
@@
where $R(M_{\rm b}=M)$ is the radial coordinate that encloses a baryonic mass at epoch $t$. The typical mass used to evaluate explodability is $M=2.5M_{\odot}$ and as close as possible to collapse. [^4] O’Connor & Ott (2011) suggested evaluating $\xi_{2.5}$ at the time of bounce, but Sukhbold & Woosley (2014) observed that similar conclusions may be obtained evaluating it at the pre-supernova link and that, indeed, the variation of the compactness after oxygen depletion is relatively minor. Following the work of Ertl et al. (2016) we also look at the behaviour of the parameters $M_{4}$ and $\mu_{4}$ defined as
@@
where $M_{r}$ is the enclosed mass inside dimensionless entropy per nucleon of $s=4k_{\rm B}$ ($k_{\rm B}$ is the Boltzmann constant). The Ertl et al. criterion consists of comparing $\mu_{4}$, which is a proxy for the mass accretion rate onto the compact remnant left after core-collapse, with the product $M_{4}\mu_{4}$, a magnitude correlated with the neutrino luminosity induced by mass accretion. This leads to a separation of models that produce neutron stars and models that produce BHs by finding which models successfully explode and which ones fail.
In Fig. 12 we show the evolution of two instances of the compactness parameter, $\xi_{M_{r}=2.5}$, $\xi_{M_{r}=M_{4}}$, where $M_{4}$ is not a fixed mass position, but its value evolves in time and is defined as in Eq. 24. A general prediction is that the larger the compactness parameter $\xi_{2.5}$, the more likely it is that the progenitor collapses forming a black hole and does not produce a successful explosion. Ugliano et al. (2012) concluded that black holes may result from stellar progenitors in which $\xi_{2.5}\gtrsim 0.35$, while models with $\xi_{2.5}\lesssim 0.15$ more likely produce neutron stars. For the range in between of these cases, $0.15\lesssim\xi_{2.5}\lesssim 0.35$, it is less certain when using only this one parameter. According to this criterion, the $25M_{\odot}$ model will collapse to a black hole employing MESA, whereas the GENEC model is more within the range of uncertainty. Noteworthy, $\xi_{2.5}$ is rising steadily prior to collapse, but reaches a value than for the MESA case. The $15M_{\odot}$ mass model, for both codes, seems much more likely to produce a neutron star and finally the $20M_{\odot}$ reaches similar values for both GENEC and MESA at the end of the evolution and lies within the uncertain range for this one parameter criterion. We notice that the compactness at the pre-SN link grows with mass, at least in the three cases at hand (though there is no monotonic dependence of $\xi_{2.5}$ on ZAMS mass; c.f., Ertl et al. 2016, Fig. 4). GENEC models with $20\,M_{\odot}$ and $25\,M_{\odot}$ show very similar values of $\xi_{2.5}$ close to collapse. To see the compactness evolution around the contracting core we also plot in Fig. 12 the quantity $\xi_{M_{r}=M_{4}}$. The location where the entropy equals $4k_{\rm B}$/baryon defines quite well the boundary of the iron core (e.g., Baron & Cooperstein 1990). We see that this compactness rises for all models during the final hours of evolution, where the value of $\xi_{M_{r}=M_{4}}$ almost doubles for all models.
@@
shown in magenta and used to analyse our lower metallicity models $20M_{\odot}$ and $25M_{\odot}$. With symbols, we plot in Fig. 13 the location of the models of Tab. 2. For the $25M_{\odot}$ cases the prediction for GENEC and MESA models are failed explosions, although, the MESA models all lie very close to the separation limit. For the $15M_{\odot}$ case, the GENEC model will most likely produce an explosion, whereas for the outcome of the MESA model remains more uncertain. As with other previously discussed aspects we see that small details and differences between the two codes could lead to an impact on the explosion success of the progenitors.
Figure 13: Position for the pre-supernova models in the ($M_{4}\cdot\mu_{4}$, $\mu_{4}$,) plane. In blue and magenta are lines separating likely explosions, below the line, and likely black hole formation, above the line. The parametric lines are taken from Ertl et al. (2016). The blue line corresponds to the calibration using the s19.8 model (Eq. (26)) and the magenta one for the w20 model (Eq. (27)).
@@ 6 Discussion
In this work, we have adapted GENEC to evolve to the pre-supernova point including a reduced nuclear network similar to the base MESA approx21_plus_co56 nuclear network.
@@
The evolution of the central density and temperature for GENEC and MESA models share almost identical slopes at the approach to collapse and in general the evolution of models shown by Fig. 5 is quite similar between codes. Looking to the central core value of the electron fraction, there is however a significant spread of the final values found between codes. The approximate network used is very dependent on the rates chosen to model the electron capture chain, and the method used to tune these rates is generally heuristic. As has been pointed out by, e.g., Renzo et al. (2024) the exact value of the profile of the electron fraction in the core is extremely important for follow-up supernovae studies and small variations here may lead to large variations in the post collapse evolution. The approximate electron capture chain is however enough to produce values at pre-collapse within the range of values found by Farmer et al. (2016) in their network and mass resolution study. Depending on the purpose of the models, the network used here may suffice. For instance, to explore further with three-dimensional studies the topology of the magnetic field and the vigour of convection in the inner regions of the star pre-collapse, this reduced network is, perhaps, the only practical one. Naturally, a detailed nucleosynthesis forecast requires the use of larger networks to also include the decrease of the electron fraction during neon and oxygen burning. This reduces the starting value of $Y_{e}$ at the beginning of silicon core burning, which can change from roughly 0.49 to $\approx$ 0.48 before even beginning the electron capture chain that we use in GeValNet25. One further point regarding networks is the use of approximate energy production. In this work, as explained in Sec. 4, for silicon burning the energy is not computed from the full network but rather by assuming QSE and thus only accounting for the energy produced by the reaction $\rm{}^{44}Ti(\alpha,\gamma)\rm^{48}Cr$ (for isotopes heavier than $\rm{}^{28}Si$). This approximation is introduced to improve the stability during the silicon core burning phase. This is of course a practical approximation to employ, but using a fully coupled network that includes the full energy calculation will undoubtedly impact the progenitor model structure by changing the burning power in convective regions. Nonetheless, small reaction networks for the calculation of the energy production rate are often used to drive the evolution as they are stable and efficient, (see for example, Weaver et al. 1978; Heger & Woosley 2010; Rauscher et al. 2002; Chieffi & Limongi 2020).
@@
We have mentioned previously the variations in the treatment in convection and overshooting varies between the two codes which can be a contributing factor to the different final results. Stellar modelling is sensitive to small variations in the size of a convective shell and it has been pointed out that understanding the small differences between codes is of utmost importance to improve stellar modelling in general (Chieffi et al. 1998). With this latest version of GENEC we aim to reduce as much as possible the differences with other stellar evolution codes given identical or very similar input physics. The EoS, opacities and network all aim to incorporate the same effects as MESA to allow for a close comparison. However, differences between codes can also be numerical and not physical in origin, e.g., the manner of the coupling of the network to the structure equations or the manner in which one implements mixing.
@@
In Tab. 2 we see that the final masses and radii are different between codes. There exist differences in the implementation of mass loss between GENEC and MESA, particularly when dealing with the treatment of super Eddington luminosity cases, which happens for the $25M_{\odot}$ model. However we show in Fig. 14 the mass loss history and during the main sequence and red giant phase the amplitude of the mass loss between codes are very similar. The difference in mass loss history rather comes from the duration of each of the different phases, where with MESA the stars tend to live longer in the main sequence and helium burning, thus leaving behind smaller total masses [^5] The exception being the $25M_{\odot}$ model that due to the treatment of mass loss being amplified in the super Eddington luminosity case loses mass faster than MESA reaching almost $10^{-4}M_{\odot}\ \rm yr^{-1}$. . The impact of mass loss is important for surface properties, such as surface abundances, or in the case of rotating stars, the surface velocities. However, for this study, where we are mainly focused on core properties and the final shell structure deep in the star, the mass loss is of much less significance. To support this argument, we refer to Farmer et al. (2016), who compared models with and without mass loss, finding only a small impact on core properties.
@@
We would finally like to highlight the variation in the Si/O shell structure between masses and codes (see Sec. 5.2). The precise nature of the region encompassing the Si-shell and the beginning of the oxygen shell can play an important role in studying supernova explosions. As pointed out by Yoshida et al. (2019), the amount of convective turbulence that a multi-D model evolved for a few minutes before collapse may produce in the silicon shell region strongly depends on the coexistence of silicon and oxygen. At the point of pre-supernova link, all models between codes and masses have very different Si/O shell structures. Some are in the process of merging, some are already merged and others well divided still. These different cases lead to convection being active in vastly different regions between models. With more detailed 1D studies and also multidimensional follow-ups of the pre-collapse evolution, we will be able to determine the origin and the importance of the differences seen in the shell structure between codes at the pre-SN link.
One final technical point that may impact the very late evolution of the core (100’s of seconds before collapse) are the inner boundary conditions. The choice of the radius of the innermost boundary is extremely important. If it is too large during silicon core burning, the convective cores can be artificially bloated by the lack of mass resolution close the the stellar center. We noted a $\sim 0.4\,M_{\odot}$ difference in the iron core mass for the $25M_{\odot}$ GENEC model when changing the inner radius from 100 km to 10 km. This highlights the importance of a follow-up mass resolution study for GENEC models, in the same vein as Farmer et al. (2016).
@@ 7 Conclusions
We have improved the GENEC code to operate in the regimes of high density and temperature characteristic of the late phases of massive stellar evolution. The improvement has included various fronts. First, we have implemented the Timmes EoS, which is thermodynamically consistent and takes into account the positron contribution to the pressure. Second, we have also extended the opacity computation to be consistent with the latest developments in the literature. Finally, we have added some elementary electron captures into the nuclear reaction network to provide a basic tracking of the neutronisation of the iron core. Our results are in good general agreement with the equivalent models run using the MESA stellar evolution code, and also with existing models in the literature employing other stellar evolution codes.
The reduction of the electron fraction in the core of the GENEC models using the GeValNet25 reaction network is consistent with the uncertainty ranges of other studies employing various network sizes (e.g., Woosley et al. 2002; Rauscher et al. 2002; Farmer et al. 2016). This reduction inevitably will lead to core collapse for the GENEC and MESA models alike, as shown by the core comparison with the Chandrasekhar mass. Thus, these models are usable to follow up on core collapse and supernova post-collapse simulations.
@@
We intend to extend upon this small grid of models to have a larger coverage of the parameter space but also include the effects of rotation and of magnetic instabilities, both of which can play crucial roles in stellar evolution and the pre-supernova model structure (e.g Meynet et al. 2004; Wheeler et al. 2015; Griffiths et al. 2022; Eggenberger et al. 2022). Additionally, the rudimentary network GeValNet25 was constructed to provide a fast and approximate network that included the reduction of the electron fraction in the core. Development of an extension of this network is ongoing to give a better estimation of the nucleosynthesis of important isotopes for observational purposes and also include more sophisticated neutrino losses during silicon core burning.

### v2

@@ 5.2 Pre-collapse structure
Figure 11: Properties of the 20 $M_{\odot}$ GENEC model before and then during a silicon-oxygen shell merger event occurring very shortly before collapse. The time gap between the two models shown is 15 s. On the left we show the convective Mach number, i.e. the ratio of the velocity predicted by mixing length theory to the sound speed, and also the energy produced by nuclear burning. On the right is shown the density and entropy structure of the shell in question.
We know that shell mergers and enhanced mixing/burning prior to core collapse can have a strong impact on the stellar structure before collapse and thus impact the supernova explosion and evolution (Müller et al. 2016; Jones et al. 2017; Rizzuti et al. 2022). We can see this also in Fig. 11. In this figure, we display the convective Mach number, entropy and density profiles of the 20 $M_{\odot}$ GENEC model before and then during a silicon and oxygen shell merger event. These profiles are taken roughly 404040\,40s before collapse, and there is a time difference of 15 s between the two models shown. We see that during the shell merger there is a rapid amplification of the energy generation by nuclear burning, an increase by more than an order of magnitude in the entire shell. This is accompanied by a steep increase in the convective mixing as the convective Mach number increases by a factor 3 in the merged shell. Furthermore, this merger event shifts the shell interface. The shift is only by $0.01M_{\odot}$; however, we note that the compactness, $\xi_{r}=\frac{M_{r}[M_{\odot}]}{(r[1000\rm km])}$, at the shell interface changes here by $8\%$ going from 0.427 to 0.464. These structural changes will impact the explosion dynamics. Therefore, knowing if the presence of these shell mergers before collapse are physical and are likely to happen is crucial. In this grid of models, shell mergers seem to be dependent on the stellar evolution code and the chosen prescriptions. These events are of course highly sensitive to differences in the network implementation and convective treatment, as such, exploration in further detail on the trigger of these shell mergers and their impact on the explosion is required.
Figure 12: Space-time diagrams, or Kippenhahn diagrams, of the $15M_{\odot}$ model for GENEC (left) and MESA (right). The grey zones show convective regions. The plot background shade displays the total energy input, the nuclear energy subtracted by the loss of thermal neutrinos in log scale. Blue represents negative energy contributions and red positive contributions . On the right axis is shown the central abundance of a given selection of species to help identify the various burning phases. The black dashed line delimits the total mass of the model.
@@ 5.3 Explodability conditions
Figure 13: Evolution prior to collapse of different instances of the compactness parameter, $\xi_{Mr=2.5}$, $\xi_{M_{r}=M_{4}}$ (see Eq. (24) for the definition.)
@@
where $R(M_{\rm b}=M)$ is the radial coordinate that encloses a baryonic mass at epoch $t$. The typical mass used to evaluate explodability is $M=2.5M_{\odot}$ and as close as possible to collapse. [^10] O’Connor & Ott (2011) suggested evaluating $\xi_{2.5}$ at the time of bounce, but Sukhbold & Woosley (2014) observed that similar conclusions may be obtained evaluating it at the pre-supernova link and that the variation of the compactness after oxygen depletion is indeed relatively minor. Following the work of Ertl et al. (2016) we also look at the behaviour of the parameters $M_{4}$ and $\mu_{4}$ defined as
@@
where $M_{r}$ is the enclosed mass inside dimensionless entropy per nucleon of $s=4k_{\rm B}$. The Ertl et al. criterion consists of comparing $\mu_{4}$, which is a proxy for the mass accretion rate onto the compact remnant left after core-collapse, with the product $M_{4}\mu_{4}$, a magnitude correlated with the neutrino luminosity induced by mass accretion. This leads to a separation of models that produce neutron stars and models that produce Black Holes by finding which models successfully explode and which ones fail.
In Fig. 13 we show the evolution of two instances of the compactness parameter, $\xi_{M_{r}=2.5}$, $\xi_{M_{r}=M_{4}}$, where $M_{4}$ is not a fixed mass position, but its value evolves in time and is defined as in Eq. 25. A general prediction is that the larger the compactness parameter $\xi_{2.5}$, the more likely it is that the progenitor collapses forming a black hole and does not produce a successful explosion. Ugliano et al. (2012) concluded that black holes may result from stellar progenitors in which $\xi_{2.5}\gtrsim 0.35$, while models with $\xi_{2.5}\lesssim 0.15$ more likely produce neutron stars. For the range in between of these cases, $0.15\lesssim\xi_{2.5}\lesssim 0.35$, it is less certain when using only this one parameter. According to this criterion, the $25M_{\odot}$ model will collapse to a black hole employing MESA, whereas the GENEC model is more within the range of uncertainty. It is noteworthy that $\xi_{2.5}$ is rising steadily prior to collapse, but reaches a lower value than the MESA model. The $15M_{\odot}$ mass model, for both codes, seems much more likely to produce a neutron star and finally the $20M_{\odot}$ reaches similar values for both GENEC and MESA at the end of the evolution and lies within the uncertain range for this one parameter criterion. We notice that the compactness at the pre-SN link grows with mass, at least in the three cases at hand (though there is no monotonic dependence of $\xi_{2.5}$ on ZAMS mass; c.f. Ertl et al. 2016, Fig. 4). GENEC models with $20\,M_{\odot}$ and $25\,M_{\odot}$ show very similar values of $\xi_{2.5}$ close to collapse. To see the compactness evolution around the contracting core we also plot in Fig. 13 the quantity $\xi_{M_{r}=M_{4}}$. The location where the entropy equals $4k_{\rm B}$/baryon defines quite well the boundary of the iron core (e.g. Baron & Cooperstein 1990). We see that this compactness rises for all models during the final hours of evolution, where the value of $\xi_{M_{r}=M_{4}}$ almost doubles for all models.
@@
shown in magenta and used to analyse our lower metallicity models $20M_{\odot}$ and $25M_{\odot}$. With symbols, we plot in Fig. 14 the location of the models of Table. 9. For the $25M_{\odot}$ cases the prediction for GENEC and MESA models are failed explosions, although, the MESA models all lie very close to the separation limit. For the $15M_{\odot}$ case, the GENEC model will most likely produce an explosion, whereas for the outcome of the MESA model remains more uncertain. As with other previously discussed aspects we see that small details and differences between the two codes could lead to an impact on the explosion success of the progenitors.
Figure 14: Position for the pre-supernova models in the ($M_{4}\cdot\mu_{4}$, $\mu_{4}$,) plane. The blue and magenta lines are separating likely explosions (below the line) and likely black hole formation (above the line). The parametric lines are taken from Ertl et al. (2016). The blue line corresponds to the calibration using the s19.8 model (Eq. (27)) and the magenta line to the w20 model (Eq. (28)).
@@ 6 Discussion
In this work, we have adapted GENEC to evolve to the pre-supernova point including a reduced nuclear network similar to the base MESA approx21_plus_co56 nuclear network. The evolution of the central density and temperature for GENEC and MESA models share almost identical slopes at the approach to collapse and in general the evolution of models shown by Fig. 6 is quite similar between codes. Looking to the central core value of the electron fraction; however, there is a significant spread of the final values found between codes. The approximate network used is very dependent on the rates chosen to model the electron capture chain, and the method used to tune these rates is generally heuristic. As has been pointed out (e.g. Renzo et al. (2024)) the exact value of the profile of the electron fraction in the core is extremely important for follow-up supernovae studies and small variations here may lead to large variations in the post-collapse evolution. The approximate electron capture chain is however enough to produce values at pre-collapse within the range of values found by Farmer et al. (2016) in their network and mass resolution study. Depending on the purpose of the models, the network used here may suffice. For instance, to explore further with three-dimensional studies the topology of the magnetic field and the vigour of convection in the inner regions of the star pre-collapse, this reduced network is, perhaps, the only practical one. Naturally, a detailed nucleosynthesis forecast requires the use of larger networks to also include the decrease in the electron fraction during neon and oxygen burning. This reduces the starting value of $Y_{e}$ at the beginning of silicon core burning, which can change from roughly 0.49 to $\approx$ 0.48 before even beginning the electron capture chain that we use in GeValNet25. One further point regarding networks is the use of approximate energy production. In this work, as explained in Sect. 4, for silicon burning the energy is not computed from the full network but rather by assuming QSE and thus only accounting for the energy produced by the reaction $\rm{}^{44}Ti(\alpha,\gamma)\rm^{48}Cr$ (for isotopes heavier than $\rm{}^{28}Si$). This approximation is introduced to improve the stability during the silicon core burning phase. This is of course a practical approximation to employ, but using a fully coupled network that includes the full energy calculation will undoubtedly impact the progenitor model structure by changing the burning power in convective regions. Nonetheless, small reaction networks for the calculation of the energy production rate are often used to drive the evolution as they are stable and efficient, (see e.g. Weaver et al. 1978; Heger & Woosley 2010; Rauscher et al. 2002; Chieffi & Limongi 2020).
@@
We have mentioned previously the variations in the treatment in convection and overshooting varies between the two codes which can be a contributing factor to the different final results. Stellar modelling is sensitive to small variations in the size of a convective shell and it has been pointed out that understanding the small differences between codes is of utmost importance to improve stellar modelling in general (Chieffi et al. 1998). With this latest version of GENEC we aim to reduce as much as possible the differences with other stellar evolution codes given identical or very similar input physics. The EoS, opacities and network all aim to incorporate the same effects as MESA to allow for a close comparison. However, differences between codes can also be numerical and not physical in origin, for example the manner of the coupling of the network to the structure equations or the manner in which one implements mixing.
@@
In Table. 9 we see that the final masses and radii are different between codes. There exist differences in the implementation of mass loss between GENEC and MESA, particularly when dealing with the treatment of super Eddington luminosity cases, which happens for the $25M_{\odot}$ model. However, we show in Fig. 15 the mass loss history and during the main sequence and red giant phase the amplitude of the mass loss between codes are very similar. The difference in mass loss history rather comes from the duration of each of the different phases, where with MESA the stars tend to live longer in the main sequence and helium burning, thus leaving behind smaller total masses [^11] The exception being the $25M_{\odot}$ model that due to the treatment of mass loss being amplified in the super Eddington luminosity case loses mass faster than MESA reaching almost $10^{-4}M_{\odot}\ \rm yr^{-1}$. . The impact of mass loss is important for surface properties, such as surface abundances, or in the case of rotating stars, the surface velocities. However, for this study, where we are mainly focused on core properties and the final shell structure deep in the star, the mass loss is of much less significance. To support this argument, we refer to Farmer et al. (2016), who compared models with and without mass loss, finding only a small impact on core properties.
@@
We would finally like to highlight the variation in the Si/O shell structure between masses and codes (see Sect. 5.2). The precise nature of the region encompassing the Si-shell and the beginning of the oxygen shell can play an important role in studying supernova explosions. As pointed out by Yoshida et al. (2019), the amount of convective turbulence that a multi-D model evolved for a few minutes before collapse may produce in the silicon shell region strongly depends on the coexistence of silicon and oxygen. At the point of pre-supernova link, all models between codes and masses have very different Si/O shell structures. Some are in the process of merging, some are already merged and others well divided still. These different cases lead to convection being active in vastly different regions between models. With more detailed 1D studies and also multidimensional follow-ups of the pre-collapse evolution, we will be able to determine the origin and the importance of the differences seen in the shell structure between codes at the pre-SN link.
One final technical point may impact the very late evolution of the core (100’s of seconds before collapse). Namely, the choice of the outer radius of the innermost mass shell (a parameter for GENEC). If it is too large during silicon core burning, the convective cores can be artificially bloated by the lack of mass resolution close the stellar centre. We noted a $\sim 0.4\,M_{\odot}$ difference in the iron core mass for the $25M_{\odot}$ GENEC model when changing the size of the innermost cell from 100 km to 10 km. This highlights the importance of a follow-up mass resolution study for GENEC models, in the same vein as Farmer et al. (2016).
@@ 7 Conclusions
We have improved the GENEC code to operate in the regimes of high-density and temperature characteristic of the late phases of massive stellar evolution. The improvements have included various fronts. First, we have implemented the Timmes EoS, which is thermodynamically consistent and takes into account the positron contribution to the pressure. Second, we have also extended the opacity computation to be consistent with the latest developments in the literature. Finally, we have added some elementary electron capture into the nuclear reaction network to provide a basic tracking of the neutronisation of the iron core. Our results are in good general agreement with the equivalent models run using the MESA stellar evolution code, and also with existing models in the literature employing other stellar evolution codes.
The reduction of the electron fraction in the core of the GENEC models using the GeValNet25 reaction network is consistent with the uncertainty ranges of other studies employing various network sizes (e.g. Woosley et al. 2002; Rauscher et al. 2002; Farmer et al. 2016). This reduction inevitably will lead to core collapse for the GENEC and MESA models alike, as shown by the core comparison with the Chandrasekhar mass. Thus, these models are usable to follow up on core collapse and supernova post-collapse simulations.
@@
We intend to extend upon this small grid of models to have a larger coverage of the parameter space, but we also plan to include the effects of rotation and of magnetic instabilities, both of which can play crucial roles in stellar evolution and the pre-supernova model structure (e.g Meynet et al. 2004; Wheeler et al. 2015; Griffiths et al. 2022; Eggenberger et al. 2022). Additionally, the rudimentary network GeValNet25 was constructed to provide a fast and approximate network that included the reduction of the electron fraction in the core. Development of an extension of this network is ongoing to give a better estimation of the nucleosynthesis of important isotopes for observational purposes and also to include more sophisticated neutrino losses during silicon core burning.